OpenStudy (anonymous):

A buoy in the harbor of San Juan, Puerto Rico, bobs up and down. The distance between the highest and lowest point is 3 feet. It moves from its highest point down to its lowest point and back to its highest point every 8 seconds. (a) Find the equation of the motion for the buoy assuming that it is at its equilibrium point at t=0 and the buoy is on its way down at that time. (b) Determine the height of the buoy at 3 seconds. (c) Determine the height of the buoy at 12 seconds.

OpenStudy (anonymous):

\[\frac{2}{3}\cos\frac{\pi}{4}\]?

OpenStudy (anonymous):

period is 16 seconds

OpenStudy (anonymous):

how?

OpenStudy (anonymous):